Show that: cos 38° cos 52° – sin 38° sin 52° = 0 kamskjhdwuhduwcnjnskjakj

Show that: 
cos 38° cos 52° – sin 38° sin 52° = 0
kamskjhdwuhduwcnjnskjakj

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

LHS= cos 38° cos 52° - sin 38° sin 52°
Simplify the given problem by converting some of the cos functions to the sin functions We know that,
cos 38° = cos (90° - 52°) = sin 52°
cos 52°= cos (90°-38°) = sin 38°
= cos (90° - 52°) cos (90°-38°) - sin 38° sin 52°
Substitute the values
= sin 52° sin 38° - sin 38° sin 52°
= 0

hence , proved

Thanks